# zkPHIRE: A Programmable Accelerator for ZKPs over HIgh-degRee, Expressive Gates

### v1

@@ VI-A3 Comparison with Prior ASICs
We now incorporate the programmable SumCheck unit into a larger accelerator that accelerates the HyperPlonk protocol, zkPHIRE. Here, our analysis focuses on yielding high-performance designs. In these comparisons, we assume the CPU runs with 32 threads, roughly 296 mm2 in 7nm. These experiments assume fixed-prime modular multipliers, which is a common technique seen in prior cryptographic accelerators and libraries [49, 9, 1].
Table II: Design Space Exploration Parameters for zkPHIRE.
@@
Figure 10: Pareto Frontiers for $2^{24}$ gates. We plot the individual Pareto curves for each bandwidth tier and the global Pareto frontier in the inset.
@@ VI-B1 Pareto Frontier Analysis
We sweep the parameters in Table II to obtain Pareto-optimal designs for each bandwidth, and then construct a global Pareto frontier across all bandwidths. Figure 10 illustrates the design space for a $2^{24}$ Jellyfish gate workload under seven bandwidth scenarios, ranging from DDR-class to HBM-scale bandwidths. We account for memory PHY overheads [49, 15, 39, 22, 47, 48, 27, 26], assuming 14.9 mm2 per HBM2 PHY and 29.6 mm2 per HBM3 PHY [2]. For $2^{24}$ Jellyfish gates, the CPU runtime is roughly 182.896 seconds. We can hit roughly $1000\times$ speedup with a 207 mm2 design and 1 TB/s off-chip bandwidth, and with $294$ mm2 and 2 TB/s, around $1400\times$ at iso-CPU area. For this high-performance regime, a Pareto design around 250 mm2 with 1 TB/s (design B) achieves nearly $2\times$ and $3\times$ speedup over 512 GB/s and 256 GB/s configurations. These results highlight a key trend in recent cryptographic accelerators where high performance demands investments in high-bandwidth memory technologies (even for small accelerators like NoCap, where the PHY cost alone is $3\times$ the compute area). That being said, with our accelerator framework, for applications with looser prover constraints, sub-DDR-scale bandwidths can still yield $100-200\times$ speedups over multi-threaded CPU baselines.
@@ VI-B2 Area, Runtime, and Power
Table III: Area and power of zkPHIRE. Other includes the PermCheck Generator, MLE Combine and SHA3 units.
@@
Table III shows an exemplar zkPHIRE design point we choose for the following analysis. In this 294 mm2 design, there are 32 MSM PEs, with 8 modular multipliers per Multifunction tree PE, and 7 EEs and 5 PLs per SumCheck PE. The design runs on 2 TB/s HBM bandwidth. We use two $\times32$ bit-sliced crossbars [47, 28, 44] in the MSM and SumCheck unit, with a shared-bus to form the on-chip bandwidth (up to 19 TB/s) between modules. Figure 12 further breaks down the runtime to compare with HyperPlonk on CPU. The colors indicate the same HyperPlonk step between CPU and zkPHIRE running at the configuration of Table III. We show the latency breakdown for Gate and Wire Identities before performing ZeroCheck masking to show their initial runtime proportions. MSMs dominate runtime before and after acceleration, though compared to zkSpeed’s CPU baseline [12], SumChecks take more runtime because the polynomials are complex, exerting more memory pressure on the CPU.
@@
VI-B3 Performance from Protocol-level Optimizations
@@
VI-B4 High-Degree Sweep
In Figure 14, we plot the runtime of the highlighted design point over custom gates represented by $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+qc$. For this family of gates, the number of witnesses is fixed, so the total MSM runtime is constant across all values of $d$. Consequently, there is a crossover point after which SumChecks dominate the total protocol runtime. This highlights a key design tradeoff: while higher-degree gates reduce overall gate count, they can potentially shift the performance bottleneck from MSMs to SumChecks. Protocol designers must therefore balance gate expressivity against the rising cost of polynomial verification.
VI-B5 Performance Comparison
Table IV compares a 300 mm2 zkPHIRE design using the same modmuls as zkSpeed+ (366 mm2 with MLE Updates pipelined into SumCheck) and without ZeroCheck masking. This is to provide as fair a comparison as possible (zkSpeed+ is roughly $10\%$ faster than zkSpeed). Speedups are reported relative to CPU. zkPHIRE is about $10\%$ slower than zkSpeed+, while offering flexibility to support arbitrary, custom gates. Table V shows zkPHIRE’s speedups at iso-CPU area (294 mm2) with fixed primes and ZeroCheck masking. We achieve $1486\times$ geomean speedup and, for the first time, bring workloads on the order of $2^{30}$ nominal gates within reach for proof generation. Lastly, in Table VI, we compare the same design (with ZeroCheck masking and fixed primes) with zkSpeed+ at iso-application, achieving geomean $11.87\times$ speedup.
Table IV: Comparison with zkSpeed+ and CPU for Vanilla gates.
@@ VI-B5 Performance Comparison
Table V: Comparison with CPU for Jellyfish gate runtime. Vanilla gate counts are shown to show original problem size.
@@
Table VI: Iso-application comparison of zkSpeed+ (Vanilla) and zkPHIRE (Jellyfish).
@@ VII Related Work
Table VII: Comparison of zkPHIRE with Prior ZKP Accelerators. N = NTT, S = SumCheck, M = MSM. Areas are in 7nm.
@@
Nocap [49] implements SumChecks using their vector architecture and chip. While SumCheck is highly parallelizable, its core operation—reducing a vector—can require repeated data access and movement on a vector processor. Reductions over a vector of length $M$ require $\log_{2}(M)$ folding steps, typically implemented as shift-and-add instructions, that must be serialized and access the vector register file. Furthermore, higher degree polynomials introduce more extensions, each of which requires a dedicated vector reduction, causing the total reduction overhead grows with degree. In contrast, zkPHIRE uses fused pipelines with tree-like, direct connections to support efficient reductions with minimal data movement. Modern ZKP protocols rely on distinct kernels like SumCheck and MSM, which differ in compute and communication patterns. zkPHIRE’s heterogeneous architecture provides specialized hardware for each, avoiding the inefficiencies of a general vector-based design balanced to handle both.
@@ VIII Conclusion
This paper presents zkPHIRE, the first accelerator to generate succinct proofs for problem sizes reaching $2^{30}$ nominal constraints. zkPHIRE supports general, programmable SumCheck, enabling ZKPs on custom, high-degree gates. zkPHIRE outperforms state-of-the-art accelerators for HyperPlonk, achieving iso-application geomean speedups of $11.87\times$ over ASIC and $1486\times$ over CPU.

### v2

@@ VI-A4 Comparison with GPU
We now incorporate the programmable SumCheck unit into a larger accelerator that accelerates the HyperPlonk protocol, zkPHIRE. Here, our analysis focuses on yielding high-performance designs. In these comparisons, we assume the CPU runs with 32 threads, roughly 296 mm2 in 7nm. These experiments assume fixed-prime modular multipliers, which is a common technique seen in prior work [54, 6, 31].
Table III: Design Space Exploration Parameters for zkPHIRE.
@@
Figure 10: Pareto Frontiers for $2^{24}$ gates. We plot the individual Pareto curves for each bandwidth tier and the global Pareto frontier in the inset. We label the highest performing design for each bandwidth tier (A - D) in the inset.
Table IV: Runtime, Area, and Bandwidth for Globally Pareto-optimal zkPHIRE Designs in Figure 10.
Design Runtime (ms) Area (mm2) BW (GB/s) CPU Speedup
[…]
@@ VI-B1 Pareto Frontier Analysis
We sweep the parameters in Table III to obtain Pareto-optimal designs for each bandwidth, then construct a global Pareto frontier across all bandwidths. Figure 10 illustrates the design space for a $2^{24}$ Jellyfish gate workload under seven bandwidth scenarios, ranging from DDR-class to HBM-scale bandwidths. We account for memory PHY overheads [54, 12, 42, 20, 52, 53, 26, 25], assuming 14.9 mm2 per HBM2 PHY and 29.6 mm2 per HBM3 PHY [18]. For $2^{24}$ Jellyfish gates, the CPU runtime is roughly 182.896 seconds. We can hit roughly $1000\times$ speedup with a 207 mm2 design and 1 TB/s off-chip bandwidth, and with $294$ mm2 and 2 TB/s, around $1400\times$ at iso-CPU area. For this high-performance regime, a Pareto design around 250 mm2 with 1 TB/s (design B) achieves nearly $2\times$ and $3\times$ speedup over 512 GB/s and 256 GB/s configurations. These results highlight a key trend in recent cryptographic accelerators where high performance demands investments in high-bandwidth memory technologies (even for small accelerators like NoCap, where the PHY cost alone is $3\times$ the compute area). That being said, with our accelerator framework, for applications with looser prover constraints, DDR-scale or lower bandwidths can still yield $100-200\times$ speedups over multi-threaded CPU baselines. In Table IV, we enumerate three additional design points; we can see that a two-order-of-magnitude speedup is achievable in under 100 mm2, relatively small by crypto-ASIC standards [28, 53, 26, 51].
@@ VI-B2 Area, Runtime, and Power
Table V: Area and power of zkPHIRE. Other includes the Permutation Quotient Generator, MLE Combine and SHA3 units.
@@
Table V shows an exemplar zkPHIRE design point we choose for the following analysis. In this 294 mm2 design, there are 32 MSM PEs, with 8 modular multipliers per Multifunction tree PE, and 7 EEs and 5 PLs per SumCheck PE. The design runs on 2 TB/s HBM bandwidth. We use two $\times32$ bit-sliced crossbars [52, 28, 48] in the MSM and SumCheck unit, with a shared-bus to form the on-chip bandwidth (up to 19 TB/s) between modules. Figure 12 further breaks down the runtime to compare with HyperPlonk on CPU. The colors indicate the same HyperPlonk step between CPU and zkPHIRE running at the configuration of Table V. We show the latency breakdown for Gate and Wire Identities before performing ZeroCheck masking to show their initial runtime proportions. MSMs dominate runtime before and after acceleration, though compared to zkSpeed’s CPU baseline [9], SumChecks take more runtime because the polynomials are complex, exerting more memory pressure on the CPU.
VI-B3 The On-Chip Memory Trade-off
Since our design space exploration (DSE) sweeps per-MLE capacity (1K - 16K words), we can assess the performance impacts of data reuse via larger scratchpads. Our DSE found that performance–area Pareto-optimal designs consistently select SumCheck units with less SRAM in favor of higher compute parallelism (e.g., more PEs/EEs/PLs). For example, from Figure 10, the Pareto front contains a 100 mm2 design (376 ms runtime) with 8 MSM PEs, 4 SumCheck PEs (4 EEs, 7 PLs), and 4K sized SRAMs. Increasing PLs from 7 to 8, we get a 104 mm2 design and 366 ms, which is still Pareto optimal. If we instead increase the SRAM size to 16K, we get a 108 mm2 design at 368 ms. While this improves performance, it is not Pareto optimal. Therefore, increasing scratchpad size does improve performance but may not offer the best performance per area.
@@ VI-B3 The On-Chip Memory Trade-off
VI-B4 Performance from Protocol-level Optimizations
@@
VI-B5 High-Degree Sweep
In Figure 14, we plot the runtime of the highlighted design point over custom gates represented by $f=q_{1}w_{1}+q_{2}w_{2}+q_{3}w_{1}^{d-1}w_{2}+qc$. For this family of gates, the number of witnesses is fixed, so the total MSM runtime is constant across all values of $d$. Consequently, there is a crossover point after which SumChecks dominate the total protocol runtime. This highlights a key design trade-off: while higher-degree gates reduce overall gate count, they can potentially shift the performance bottleneck from MSMs to SumChecks. Protocol designers must therefore balance gate expressivity with the rising cost of polynomial verification.
Table VI: Runtime Comparison with zkSpeed+ and CPU for Vanilla gates (speedups are relative to CPU).
@@ VI-B5 High-Degree Sweep
VI-B6 Performance Comparison
Table VI compares a 300 mm2 zkPHIRE design using the same modular multipliers as zkSpeed+ (366 mm2 with MLE Updates pipelined into SumCheck) and without ZeroCheck masking. This is to provide as fair a comparison as possible (zkSpeed+ is roughly $10\%$ faster than zkSpeed). Speedups are reported relative to CPU. zkPHIRE is about $10\%$ slower than zkSpeed+, while offering flexibility to support arbitrary, custom gates. Table VII shows zkPHIRE’s speedups at iso-CPU area (294 mm2) with fixed primes and ZeroCheck masking, achieving $1486\times$ geomean speedup and generating proofs for workloads upwards of $2^{30}$ nominal gates. Lastly, Table VIII compares the same design (with ZeroCheck masking and fixed primes) with zkSpeed+ at iso-application, achieving geomean $11.87\times$ speedup.
Table VII: Runtime and speedup compared to CPU for Jellyfish gates. Vanilla gate counts are shown to show original problem size.
@@ VI-B6 Performance Comparison
Table VIII: Iso-application comparison of zkSpeed+ (Vanilla)
and zkPHIRE (Jellyfish).
@@ VII Related Work
Table IX: Comparison of zkPHIRE with Prior ZKP Accelerators.
N = NTT, S = SumCheck, M = MSM. Areas are in 7nm.
@@
Prior work on cryptographic acceleration has largely focused on Multi-Party Computation and Homomorphic Encryption [28, 25, 56, 42, 15, 11, 45, 41, 26, 22, 12, 27]. Research on ZKP acceleration is relatively newer, and has focused primarily on MSMs and NTTs [32, 23, 36, 74, 68, 63, 73, 38, 34, 50, 7, 71, 72, 75, 62, 33, 19, 59]. The community has broadened its focus to include other ZKP kernels such as the SumCheck protocol and Merkle tree-based commitments. Prior work has explored accelerating these kernels on GPUs [35] and ASICs [54, 9]. We compare zkPHIRE with three ASICs that accelerate proof-generation end-to-end: SZKP, NoCap, and zkSpeed.
Protocol Acceleration: SZKP is a Groth16 [17] accelerator that speeds up all MSMs, including sparse G2 MSMs, achieving a geomean speedup of 493× over CPU. While a direct comparison is not apples-to-apples due to differing proof sizes and trusted setup assumptions, Groth16 and HyperPlonk target similar application domains requiring small, efficiently verifiable proofs. Unlike Groth16, HyperPlonk supports a universal setup and avoids an expensive trusted setup ceremony [64]. Following zkSpeed, we report results for a scaled SZKP+ configuration that adopts more efficient MSMs and the BLS12-381 curve. NoCap accelerates Spartan+Orion proofs using a vector architecture. It relies on Merkle tree-based commitments, resulting in larger proofs and higher verifier costs, and is better suited for settings where proof succinctness is less critical or number of verifiers is small. zkSpeed is, to our knowledge, the only prior accelerator targeting HyperPlonk end-to-end, achieving a geometric mean speedup of 801× over CPU. For our comparison here, we use zkSpeed+ as we have done earlier.
@@
Limitations of prior work: Prior work [54] accelerates SumCheck using vector processors, but vector-style reductions introduce significant overhead: reductions over length-$M$ vectors require $\log_{2}(M)$ serialized folding steps, repeated register-file accesses, and additional reductions for higher-degree extensions. zkPHIRE avoids these bottlenecks by using fused, tree-structured reduction pipelines with minimal data movement. While prior work [54, 68] can scale to workloads of $2^{30}$ constraints, their reliance on NTT-based matrix encodings introduces substantial latency overhead, e.g., off-chip transposes and recursive decompositions. In contrast, zkPHIRE has straightforward memory accesses, and given the streaming architecture, can scale well beyond $2^{30}$ constraints, as long as the working set fits in off-chip memory. For ZKML or other problems of this scale, zkPHIRE would need better data orchestration schemes (e.g. fetching from the cloud). We leave this for future work.
@@ VIII Conclusion
This paper presents zkPHIRE, the first accelerator to support programmable SumCheck for succinct proof generation at very large scales, enabling efficient ZKPs over custom, high-degree gates. On HyperPlonk baselines, zkPHIRE achieves iso-application geomean speedups of $11.87\times$ over prior ASICs and $1486\times$ over CPU. Its design abstracts SumCheck into common primitives—updates, extensions, and products—and exploits parallelism within tables, across tables, and across evaluation products to scale across diverse polynomial structures. By streaming polynomial tables through compact scratchpads and coupling compute units via a Multifunction Forest, zkPHIRE improves resource sharing and reduces SRAM and datapath overheads. Finally, a systematic design-space exploration balances parallelism, area efficiency, and performance to support a wide range of polynomials and custom gate constructions.
Acknowledgements
This work was supported in part by DARPA’s Hybrid Electromagnetic Side-Channel and Interactive-Proof Methods to Detect and Amend LogicaL Rifts (HEIMDALLR) program (Grant HR0011-25-C-0300), NSF CAREER Award #2340137, an NSF RINGS Award, and a gift award from Google.